The arc length of a curve described by r(t)=⟨cos(3t),sin(3t),3t⟩ for t∈[0,π] is
Select one:
a. 3√2π

b. 3√3π

c. √3π

d. 6√2π

e. √2π

f. 2√2π
1
Expert's answer
2020-05-19T08:34:38-0400

"r(t)=[{\cos(3t),\sin(3t),3t}]"

"r'(t)=[{-3\sin(3t),3\cos(3t),3}]"

"\begin{vmatrix}
 r'(t)
\end{vmatrix}=\sqrt{(-3\sin(3t)^2+(3\cos(3t))^2+3^2}=\\=3\sqrt{2}"

The length of the curve ="\int_0^\pi3\sqrt{2}dt=3\sqrt{2}\pi"

Answer: a."3\sqrt2\pi" .